# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ Abstract
The pursuit of human-level artificial intelligence (AI) has significantly advanced the development of autonomous agents and Large Language Models (LLMs). LLMs are now widely utilized as decision-making agents for their ability to interpret instructions, manage sequential tasks, and adapt through feedback. This review examines recent developments in employing LLMs as autonomous agents and tool users and comprises seven research questions. We only used the papers published between 2023 and 2025 in conferences of the A* and A rank and Q1 journals. A structured analysis of the LLM agents’ architectural design principles, dividing their applications into single-agent and multi-agent systems, and strategies for integrating external tools is presented. In addition, the cognitive mechanisms of LLM, including reasoning, planning, and memory, and the impact of prompting methods and fine-tuning procedures on agent performance are also investigated. Furthermore, we evaluated current benchmarks and assessment protocols and have provided an analysis of 68 publicly available datasets to assess the performance of LLM-based agents in various tasks. In conducting this review, we have identified critical findings on verifiable reasoning of LLMs, the capacity for self-improvement, and the personalization of LLM-based agents. Finally, we have discussed ten future research directions to overcome these gaps.
@@ 2 Related works
This section comprehensively analyzes the existing survey literature on LLMs as agents and tool users. The emergence of LLMs as autonomous agents and tools has sparked interest in AI research. For example, Ferrag et al. [27] provided a basic taxonomy of agents based on LLM, describing reasoning, planning, and tool use capabilities. The survey cataloged over 60 benchmarks and systematically reviewed frameworks for agentic behavior. Li et al. [28] also offered an analysis of three agentic paradigms: tool use, retrieval-based planning, and feedback-driven learning. They categorized LLM agent roles, discussed the limitations of task-agnostic frameworks, and proposed directions for composable and generalizable agent development. Similarly, Xu et al. [29] focused on tool-augmented LLMs and outlined strategies for integrating external functionalities, including prompting, multimodal interaction, and agent coordination.
@@
Table 1 demonstrates a comparative analysis of seven prominent surveys against our review content, structured around the research questions (RQs): baseline LLMs used (RQ1), integration of external tools (RQ2), frameworks for building LLM agents (RQ3), reasoning, planning, and memory capabilities (RQ4), prompting and fine-tuning strategies (RQ5), evaluation and benchmarks (RQ6) and concerns and limitations (RQ7).
Existing studies often emphasize the discussion of integrating external tools across LLM agent workflows (i.e., RQ2); however, foundational dimensions, such as baseline LLM usage and the impact of prompting, fine-tuning, and memory enhancement, receive comparatively limited attention. In contrast, by addressing all seven key areas, we propose our own taxonomy of agentic systems based on LLM, presented in Table 2. This taxonomy extends existing frameworks by organizing the field into core methodologies, agent capabilities, domain-specific applications, evaluation and safety aspects, and human-agent interaction. Our holistic approach distinguishes us as the most comprehensive to date, addressing the foundational and emergent dimensions of agents and tools based on LLM, offering a unified perspective on their architectures, capabilities, and future directions.
@@ 3 Methodology
This study uses a clear and organized methodology to explore the evolving field of LLM agents. The analysis is guided by targeted RQs that aim to clarify the basic structures, capabilities, and environments of these agents. The literature selection process included a wide range of studies that focus on foundational structures, new methods, and practical implementation approaches. The selected works were organized to allow for a detailed look at new trends, system designs, and mechanisms that allow agent-like behaviors in LLMs.
@@ 3.2 Search strategies
Search sources. The literature search focused on peer-reviewed publications from high-impact journals and conference proceedings in AI, machine learning, and natural language processing. The main sources included NeurIPS, ICML, ICLR, ACL, EMNLP, AAAI, EAAI, CVPR, ICCV, ACM, Nature Machine Intelligence, NPJ Digital Medicine, ACM Transaction, IEEE Transaction and AI journals. These venues were selected for their reliable coverage of recent advances in LLMs and agent-based systems. The search looked for publications from 2023 to the present to capture the latest developments in this fast-moving field.
Search terms. A focused set of search terms was created to target the overlap of LLM and autonomous agents. These included: “Large Language Model agents,” “LLM-based agents,” “multi-agent LLM systems,” “tool-augmented LLMs,” “LLM planning and reasoning,” “LLM self-reflection,” “autonomous LLM agents,” “communicative LLM agents,” “LLM agent frameworks,” “embodied LLM agents,” and “LLM tool integration.” The terms were gradually refined on the basis of keywords found in key works and emerging trends in the literature.
@@ 3.3 Selection criteria
Provider Model Variants Applications Studies
OpenAI GPT-4, GPT-3.5, GPT-2, Text-DaVinci-003, Code-DaVinci-002, o1-preview, o1-mini Core reasoning, planning, benchmark reference, multi-agent collaboration, embodied agents [108, 71, 109, 98, 86, 90, 110, 40, 104, 38, 39, 111, 69, 34, 35, 42, 112, 19, 84, 102, 101, 53, 47, 92, 95, 113, 48, 114, 80, 88, 87, 115, 116, 64, 85, 37, 63, 103, 117, 118, 100, 94, 119, 120, 121, 78, 70, 77, 56, 76, 122, 68, 123, 89]
Anthropic Claude 1, 2, 3 Opus, 3.5, 3.7 Sonnet, 3 Haiku Benchmarking, time series learning, hallucination detection, R&D comparison, behavior elicitation [41, 90, 61, 94, 51, 105, 89, 124, 99, 19, 110, 116]
Google DeepMind Gemini 1.0 Pro, Gemini 1.5 Pro, Gemini Pro, PaLM 2, FLAN-T5 Agent development, distraction reasoning, comparative evaluation, enterprise agent foundation [94, 83, 91, 109, 39, 107, 101, 105, 41, 125]
Others Mistral-medium-2312, Chinchilla-LM (70B), HyperCLOVA (82B) Benchmark tasks, collaborative agents, large-scale agent frameworks [105, 126, 81]
@@ 3.4 Article selection
We systematically reviewed agentic LLM systems and the role of integrated tools within such systems. Articles were selected from A* and A-ranked conferences and the leading Q1 journals. Figure 3 presents two complementary views of the publication distribution: (A) the distribution of selected articles between conferences and journals, and (B) the monthly distribution of published articles from 2023 to 2025. The review considered studies published between 2023 and 2025, with the majority published in 2024 and 2025. A total of 108 articles were included based on predefined IC, and 75 were published in A* conferences, highlighting our concentration of high impact research. Furthermore, Table 3 summarizes the geographical distribution of the publications, indicating that China and the United States account for the majority of contributions.
4 Baseline Large Language Models for agent frameworks
@@ 4.1 Proprietary LLMs for agentic applications
Among the available language models, the Generative Pre-trained Transformer (GPT) family, most notably GPT-4 and its variants, remains the most widely adopted foundation for agent implementations [108]. These models often serve as the core decision-making component or serve as reference baselines to assess the relative performance of novel techniques. For example, GPT models such as GPT-4 and GPT-3.5 provide a consistent point of comparison when evaluating the effectiveness of other open source LLMs [108, 71, 109, 98, 86, 90, 110]. In certain studies, GPT models have been integrated with other open-source counterparts to function collaboratively within multi-agent settings [40, 104, 38, 39, 111, 69]. In addition to being used for benchmarking and integrating other models, GPT variants are frequently used as foundational models in several agent-based systems explored in studies [40, 71, 34, 35, 42, 112, 104, 109, 19, 84, 38, 102, 101, 53, 47, 39, 92, 95, 113, 48, 114, 80, 88, 87, 115, 116, 64, 85, 37, 63, 103, 117, 118, 64, 100, 94, 119, 120, 121, 69, 78, 70, 77, 56, 76]. Based on our analysis, we identified approximately 55 studies that employed GPT-4 or its variants and around 23 studies that used GPT-3.5. Beyond GPT-4 and GPT-3.5, our review also identified the use of additional OpenAI models, including Code-DaVinci-002 [122], Text-DaVinci-003 [118, 122, 68, 123], and GPT-2 [42] as agents. Furthermore, Wijk et al. used the variant o1-preview [89], and o1-mini appeared in the study by Arumugam et al. [115].
The Claude model family, developed by Anthropic, has emerged as key players in the proprietary LLM domain and is frequently regarded as a primary competitor to OpenAI models [41, 90, 61, 94]. The Claude 3 lineup includes the Claude 3 Opus, the most advanced and capable model in the family, which has been used in several studies [51, 105]. Claude 3.5 and 3.7 Sonnet, recognized for their improved speed and capability, have been used as agents in several tasks such as LLM behavior elicitation, time-series ML engineering, and comparing research and development (R&D) capabilities [89, 124, 99]. Earlier generation models, such as Claude-2, have been examined in multiple comparative evaluation studies [19, 110], and the research capabilities of agents powered by Claude-1 have been benchmarked [116].
@@
Model Family Model Variants Applications Studies
Meta LLaMA LLaMA 2 (7B, 13B, 70B), LLaMA 3 (8B, 70B), LLaMA-3.1 (11B), LLaMA-3.2 (90B), CodeLLaMA (7B, 34B), ToolLLaMA-7B, COLLaMA-2 Zero-shot eval, LoRA fine-tuning, tool use, programming agents, collaborative agents [60, 58, 74, 129, 86, 71, 52, 42, 45, 102, 57, 87, 130, 93, 73, 113, 46, 72, 131, 132, 79, 53, 50, 90, 95, 55, 36, 109]
Mistral AI Mistral-7B, Mixtral-8×7B Efficient inference, strong generation quality, GPT-3.5 comparable coding [71, 91, 66, 97, 49, 50, 133, 105]
Google Gemma Gemma-7B, Gemma-2 (2B) Retrieval-augmented planning, low-resource numerical agents [44, 50, 134]
Alibaba Qwen Qwen-max, Qwen 2.5 (72B) Star-Agents, SMAC reinforcement [108, 44]
DeepSeek AI DeepSeek-r1-70B, DeepSeek-7B, DeepSeek V2.5 Top-tier open-source agent models [108, 75, 49]
Vicuna/WizardLM Vicuna-13B, Wizard-Vicuna-30B, WizardLM-70B Conversational agents, high-capacity chat models [91, 98, 54, 110, 129, 135, 136, 72]
Zhipu AI GLM-4, ChatGLM3 General chat and task execution [44]
Microsoft Phi-2, PHI-3.5 Mini Compact inference agents [108]
OpenChat/Baichuan OpenChat-3.5, Baichuan-13B Dialogue-optimized agents [110]
LongChat LongChat-7B Long-context critic agents [111]
Multi-modal LLaVA-7B, LLaVA-v1.6-mistral-7B, CogVLM-17B Vision-language agents, GUI-based interactions [133, 62, 59]
Generative/Visual CLIP, Stable Diffusion XL, VQ-GAN Image generation, visual embedding for agents [58]
In LLM agent research, Google’s proprietary models are frequently featured, particularly in frameworks involving multiple model evaluations. The Gemini series is the most prominent [94, 83], with Gemini Pro emerging as the version most often used in agent development and evaluation [91, 109, 39, 107]. Gemini 1.5 Pro, known for its large context window, has been used in LLM-based agent reasoning under distraction tasks [101], and Gemini 1.0 Pro can be found in specific LLM agent benchmark studies [105, 41]. The PaLM lineup is also represented, and PaLM 2 appears in the work of Zhang et al. [41]. Other significant models include FLAN-T5, which is noted as a foundational component for enterprise-level agents [125].
In addition to proprietary models from well-known tech companies, the Mistral medium-2312 [105], Chinchilla-LM (70B) [126], and HyperCLOVA (82B) [81] models are also included in our review of LLM agents.

### v2

@@ Abstract
The pursuit of human-level artificial intelligence (AI) has significantly advanced the development of autonomous agents and Large Language Models (LLMs). LLMs are now widely utilized as decision-making agents for their ability to interpret instructions, manage sequential tasks, and adapt through feedback. This review examines recent developments in employing LLMs as autonomous agents and tool users and comprises seven research questions. We only used the papers published between 2023 and 2025 in conferences of the A* and A-ranked and Q1 journals. A structured analysis of the LLM agents’ architectural design principles, dividing their applications into single-agent and multi-agent systems, and strategies for integrating external tools is presented. In addition, the cognitive mechanisms of LLMs, including reasoning, planning, and memory, and the impact of prompting methods and fine-tuning procedures on agent performance are also investigated. Furthermore, we have evaluated current benchmarks and assessment protocols and provided an analysis of 68 publicly available datasets to assess the performance of LLM-based agents in various tasks. In conducting this review, we have identified critical findings on verifiable reasoning of LLMs, the capacity for self-improvement, and the personalization of LLM-based agents. Finally, we have discussed ten future research directions to overcome these gaps.
@@ 2 Related works
This section comprehensively analyzes the existing survey literature on LLMs as agents and tool users. The emergence of LLMs as autonomous agents and tool users has sparked interest in AI research. For example, Ferrag et al. [26] provided a basic taxonomy of agents based on LLM, describing reasoning, planning, and tool use capabilities. Their survey cataloged over 60 benchmarks and systematically reviewed frameworks for agent behavior. Li et al. [27] also offered an analysis of three agentic paradigms: tool use, retrieval-based planning, and feedback-driven learning. They categorized LLM agent roles, discussed the limitations of task-agnostic frameworks, and proposed directions for composable and generalizable agent development. Similarly, Xu et al. [28] focused on tool-augmented LLMs and outlined strategies for integrating external functionalities, including prompting, multimodal interaction, and agent coordination.
@@
Table 1 presents a comparative analysis of seven prominent surveys against our review content, structured around the research questions (RQs): baseline LLMs used (RQ1), integration of external tools (RQ2), frameworks for building LLM agents (RQ3), reasoning, planning, and memory capabilities (RQ4), prompting and fine-tuning strategies (RQ5), evaluation and benchmarks (RQ6) and concerns and limitations (RQ7).
Existing studies often emphasize the discussion of integrating external tools across LLM agent workflows (i.e., RQ2); however, foundational dimensions, such as baseline LLM usage and the impact of prompting, fine-tuning, and memory enhancement, receive comparatively limited attention. In contrast, addressing all seven key areas, we propose our own taxonomy of agentic systems based on LLM, presented in Table 2. This taxonomy extends existing frameworks by organizing the field into core methodologies, agent capabilities, domain-specific applications, evaluation and safety aspects, and human-agent interaction. Our holistic approach distinguishes us as the most comprehensive to date, addressing the foundational and emergent dimensions of agents and tools based on LLM, offering a unified perspective on their architectures, capabilities, and future directions.
@@ 3 Methodology
This study uses a clear and organized methodology to explore the evolving field of LLM agents. The analysis is guided by targeted RQs that aim to clarify the basic structures, capabilities, and environments of these agents. The literature selection process included a wide range of studies that focused on foundational structures, new methods, and practical implementation approaches. The selected works were organized to allow for a detailed look at new trends, system designs, and mechanisms that allow agent-like behaviors in LLMs.
@@ 3.2 Search strategies
Search sources. The literature search focused on peer-reviewed publications from high-impact journals and conference proceedings in AI, machine learning, and natural language processing. The main sources included NeurIPS, ICML, ICLR, ACL, EMNLP, AAAI, EAAI, CVPR, ICCV, ACM, Nature Machine Intelligence, NPJ Digital Medicine, ACM Transactions, IEEE Transactions, and AI journals. These venues were selected for their reliable coverage of recent advances in LLMs and agent-based systems. The search looked for publications from 2023 to the present to capture the latest developments in this fast-moving field.
Search terms. A focused set of search terms was created to target the overlap of LLM and autonomous agents. These included: “Large Language Model agents,” “LLM-based agents,” “multi-agent LLM systems,” “tool-augmented LLMs,” “LLM planning and reasoning,” “LLM self-reflection,” “autonomous LLM agents,” “communicative LLM agents,” “LLM agent frameworks,” “embodied LLM agents,” and “LLM tool integration.” The terms were gradually refined on the basis of keywords found in the literature and emerging trends in the literature.
@@ 3.3 Selection criteria
Provider Model Variants Primary Applications Key Differences
OpenAI [104, 68, 105, 94, 83, 23, 106, 39, 100, 37, 38, 107, 66, 33, 34, 41, 108, 19, 81, 98, 97, 51, 45, 88, 91, 109, 46, 110, 77, 85, 84, 111, 112, 62, 82, 36, 61, 99, 113, 114, 96, 90, 115, 116, 117, 75, 67, 74, 54, 73, 118, 65, 119, 86] GPT-4, GPT-3.5, GPT-2, Text-DaVinci-003, Code-DaVinci-002, o1-preview, o1-mini Core reasoning, planning, benchmark reference, multi-agent collaboration, embodied agents multimodal, tool-use/agents, advanced reasoning
Anthropic [40, 23, 59, 90, 49, 101, 86, 120, 95, 19, 106, 112] Claude 1, 2, 3 Opus, 3.5, 3.7 Sonnet, 3 Haiku Benchmarking, time series learning, hallucination detection, R&D comparison, behavior elicitation ultra-long context, safety-first alignment, strong instruction following
Google DeepMind [90, 80, 87, 105, 38, 103, 97, 101, 40, 121] Gemini 1.0 Pro, Gemini 1.5 Pro, Gemini Pro, PaLM 2, FLAN-T5 Agent development, distraction reasoning, comparative evaluation, enterprise agent foundation native multimodal, million-token context, Google-ecosystem integration
Others [101, 122, 78] Mistral-medium-2312, Chinchilla-LM (70B), HyperCLOVA (82B) Benchmark tasks, collaborative agents, large-scale agent frameworks inference-efficient small models, data-efficient scaling
@@ 3.4 Article selection
We systematically reviewed agentic LLM systems and the role of integrated tools within such systems. Articles were selected from A* and A-ranked conferences and leading Q1 journals. Figure 3 presents two complementary views of the publication distribution: (A) the distribution of selected articles between conferences and journals, and (B) the monthly distribution of published articles from 2023 to 2025. The review considered studies published between 2023 and 2025, with the majority appearing in 2024 and 2025. A total of 108 articles were included based on predefined inclusion (IC) and exclusion (EC) criteria. Among those, 75 were published in A* conferences, highlighting our concentration of high-impact research. Furthermore, Table 3 summarizes the geographical distribution of the publications, indicating that China and the United States account for the majority of contributions.
3.5 Thematic scoring for article selection
A total of 108 papers were included in this review. To ensure a systematic and objective selection, each paper was evaluated using a structured thematic scoring framework, which quantified coverage across eight predefined dimensions:
  • Focus on LLM Agentic Behavior
  • Training Mechanisms for Autonomy
  • Interaction with External Tools or Environments
  • Frameworks for Single/Multi-Agent Systems
  • Reasoning, Planning, Memory, or Self-Reflection
  • Prompting, Fine-Tuning, or Memory Augmentation
  • Evaluation Metrics and Benchmarks for Agentic Performance
  • Challenges, Limitations, and Ethical Considerations
Each dimension was scored on a scale from 0, 1, or 2 per co-author, where (0) means "no", (1) means "partially", and (2) means "yes". The score for each question is a cumulative score of four co-authors, with a maximum score of 8 for each question (i.e., if co-author one gives 2, co-author two gives 1, co-author three gives 0, and co-author four gives 2, then the thematic score of this question is 5). Then the total score for each paper was computed as the sum across all eight dimensions, with a maximum possible score of 64. Papers with higher total scores were considered more comprehensive, while all included papers met a minimum inclusion threshold (total score $\geq$ 48).
The scoring was conducted independently by four co-authors with expertise in LLMs and autonomous systems. Any disagreements in scoring were resolved through discussion until consensus was reached, ensuring high reliability and reproducibility. This structured approach allows for transparent, objective synthesis of the literature, highlighting methodological trends, key contributions, and gaps in the study of LLMs as autonomous agents.
The complete thematic scoring for all 108 papers is provided as supplementary material in Table S1.111https://github.com/mak-raiaan/LLMAgentsReview
4 Baseline LLMs for agent frameworks
@@ 4.1 Proprietary LLMs for agentic applications
Among the available language models, the Generative Pre-trained Transformer (GPT) family, most notably GPT-4 and its variants, remains the most widely adopted foundation for agent implementations [104]. These models often serve as the core decision-making component or serve as reference baselines to assess the relative performance of novel techniques. For example, GPT models such as GPT-4 and GPT-3.5 provide a consistent point of comparison when evaluating the effectiveness of other open source LLMs [104, 68, 105, 94, 83, 23, 106]. In certain studies, GPT models have been integrated with other open-source counterparts to work collaboratively within multi-agent settings [39, 100, 37, 38, 107, 66]. In addition to being used to benchmark and integrate other models, GPT variants are frequently used as foundational models in several agent-based systems explored in studies [39, 68, 33, 34, 41, 108, 100, 105, 19, 81, 37, 98, 97, 51, 45, 38, 88, 91, 109, 46, 110, 77, 85, 84, 111, 112, 62, 82, 36, 61, 99, 113, 114, 62, 96, 90, 115, 116, 117, 66, 75, 67, 74, 54, 73]. Based on our analysis, we identified approximately 55 studies that employed GPT-4 or its variants and around 23 studies that used GPT-3.5. Beyond GPT-4 and GPT-3.5, our review also identified the use of additional OpenAI models, including Code-DaVinci-002 [118], Text-DaVinci-003 [114, 118, 65, 119], and GPT-2 [41] as agents. Furthermore, Wijk et al. used the variant o1-preview [86], and o1-mini appeared in the study by Arumugam et al. [111].
The Claude model family, developed by Anthropic, has emerged as a key player in the proprietary LLM domain and is often considered as a primary competitor to OpenAI models [40, 23, 59, 90]. The Claude 3 lineup includes the Claude 3 Opus, the most advanced and capable model in the family, which has been used in several studies [49, 101]. Claude 3.5 and 3.7 Sonnet, recognized for their improved speed and capability, have been used as agents in several tasks such as LLM behavior elicitation, time-series ML engineering, and comparing research and development (R&D) capabilities [86, 120, 95]. Earlier generation models, such as Claude-2, have been examined in multiple comparative evaluation studies [19, 106], and the research capabilities of agents powered by Claude-1 have been benchmarked [112]. Moreover, GPT models excel in reasoning, instruction-following, and tool integration, making them highly versatile and widely adopted for general-purpose and multi-agent applications. However, Claude 3 models, especially Opus and 3.5/3.7 Sonnet, provide faster responses and stronger performance in R&D.
@@
Model Family Model Variants Primary Applications Key Differences
Meta LLaMA [58, 56, 71, 124, 83, 68, 50, 41, 43, 98, 55, 84, 125, 89, 70, 109, 44, 69, 126, 64, 76, 51, 48, 23, 91, 53, 35, 105] LLaMA 2 (7B, 13B, 70B), LLaMA 3 (8B, 70B), LLaMA-3.1 (11B), LLaMA-3.2 (90B), CodeLLaMA (7B, 34B), ToolLLaMA-7B, COLLaMA-2 Zero-shot eval, LoRA fine-tuning, tool use, programming agents, collaborative agents Open-weights, broad sizes, strong coding, huge ecosystem, multilingual, tool-use
Mistral AI [68, 87, 63, 93, 47, 48, 127, 101] Mistral-7B, Mixtral-8×7B Efficient inference, strong generation quality, GPT-3.5 comparable coding Inference-efficient, Mixture of Experts (MoE), fast throughput, long-context, strong coding
Google Gemma [42, 48, 128] Gemma-7B, Gemma-2 (2B) Retrieval-augmented planning, low-resource numerical agents Lightweight, safety-tuned, efficient fine-tuning
Alibaba Qwen [104, 42] Qwen-max, Qwen 2.5 (72B) Star-Agents, SMAC reinforcement Chinese-English, long-context, tool-use & enterprise features, multimodal
DeepSeek AI [104, 72, 47] DeepSeek-r1-70B, DeepSeek-7B, DeepSeek V2.5 Open-source agent models Reasoning-focused, math/code strength, cost-efficient, RL-style alignment, long-context
Vicuna/WizardLM [87, 94, 52, 106, 124, 129, 130, 69] Vicuna-13B, Wizard-Vicuna-30B, WizardLM-70B Conversational agents, high-capacity chat models Instruction-tuned chat, easy fine-tune
Zhipu AI [42] GLM-4, ChatGLM3 General chat and task execution Chinese-first bilingual, tool-use & agents, long-context, enterprise stack
Microsoft [104] Phi-2, Phi-3.5 Mini Compact inference agents Ultra-small, on-device ready, textbook-style data curation, inference-efficient
OpenChat/Baichuan [106] OpenChat-3.5, Baichuan-13B Dialogue-optimized agents Conversational alignment, Chinese-English, open-weights
LongChat [107] LongChat-7B Long-context critic agents Extra-long context, memory retention, position scaling tricks
Multimodal [127, 60, 57] LLaVA-7B, LLaVA-v1.6-mistral-7B, CogVLM-17B Vision-language agents, GUI-based interactions Vision-language, image grounding, visual question-answering & captioning, perceptual reasoning
Generative/Visual [56] CLIP, Stable Diffusion XL, VQ-GAN Image generation, visual embedding for agents Image generation, latent diffusion, visual embeddings
In LLM agent research, Google’s proprietary models are frequently featured, particularly in frameworks involving multiple model evaluations. The Gemini series is the most prominent [90, 80], with Gemini Pro emerging as the version most often used in agent development and evaluation [87, 105, 38, 103]. Gemini 1.5 Pro, known for its large context window, has been used in LLM-based agent reasoning under distraction tasks [97], and Gemini 1.0 Pro can be found in specific LLM agent benchmark studies [101, 40]. The PaLM lineup is also represented, and PaLM 2 appears in the work of Zhang et al. [40]. Other significant models include FLAN-T5, which is noted as a foundational component for enterprise-level agents [121]. In addition to proprietary models from well-known tech companies, the Mistral medium-2312 [101], Chinchilla-LM (70B) [122], and HyperCLOVA (82B) [78] models are also included in our review of LLM agents.